OpenStudy (errinss):

radical square root, can anyone help me understand it a little more? will fan and medal

OpenStudy (errinss):

@jim_thompson5910 @A_Burning_Masquerade

jimthompson5910 (jim_thompson5910):

what's your question about radicals and square roots?

OpenStudy (errinss):

i dont really understand how to solve radicals, you-tube doesn't help

jimthompson5910 (jim_thompson5910):

which equation are you working with?

OpenStudy (errinss):

that is a radical right?

jimthompson5910 (jim_thompson5910):

yes a square root is a form of a radical

OpenStudy (errinss):

heh thought so XP

jimthompson5910 (jim_thompson5910):

a perfect square is a number where it represents areas of squares with whole numbered side lengths example: if a square has side length of 3, then the area is 9. The number 9 is a perfect square. 49 is a perfect square because 7^2 = 7*7 = 49 the list of perfect squares are... 1,4,9,16,25,36,49, etc etc the list goes on forever

jimthompson5910 (jim_thompson5910):

which numbers in this list 1,4,9,16,25,36,49 are factors of 54?

OpenStudy (errinss):

9

jimthompson5910 (jim_thompson5910):

we can break up 54 into 9*6 meaning that \[\Large \sqrt{54} = \sqrt{9*6}\] agreed?

OpenStudy (errinss):

yea, i conquer

jimthompson5910 (jim_thompson5910):

we now use this rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\] to break up the root

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\] \[\Large \sqrt{9*6} = \sqrt{9}*\sqrt{6}\] \[\Large \sqrt{9*6} = 3*\sqrt{6}\]

jimthompson5910 (jim_thompson5910):

Take note how the square root of 9 is 3 this is because 3 squared = 3^2 = 3*3 = 9. In a sense, we're going backwards from squaring

OpenStudy (errinss):